OpenStudy (anonymous):

find the volume of a circular cone if the diameter of the base is 12 inches and the height is 2 feet?

OpenStudy (anonymous):

\[(\pi \times r^2 \times h)/3\]

OpenStudy (anonymous):

@star tats nt the formula

OpenStudy (anonymous):

soo can i take 12X12=144

OpenStudy (anonymous):

x2

OpenStudy (anonymous):

=288pi in3

OpenStudy (anonymous):

@star tats nt the formula This being the second dumb mistake I've made tonight, perhaps a sign to go to sleep, yes? ha.

OpenStudy (anonymous):

its ok. haha. it happens. just gotta big test tmrw evening and i wanna make sure i am on the right track

OpenStudy (anonymous):

You might try plugging in for: r = 6, h = 2...and your answer may magically reveal 24pi !

OpenStudy (anonymous):

so am i over thinking this than?

OpenStudy (anonymous):

6x6=36 x2=72/3

OpenStudy (anonymous):

=24?

OpenStudy (anonymous):

Remember to plug in 6, not 12 , for the radius into the formula that savinand gave earlier and you should have 24pi in^3.

OpenStudy (anonymous):

ok because the diameter of the base is already given..you use the radius of the base to find the volume. okay. thank u. i was given a practice sheet either from a copy made from someone elses answers but i wanted to practice myself and i had a feeling 288 was wrong